OpenStudy (anonymous):

Select True or False. The expression 3x − 4(2x − 5) is equivalent to the expression 20 − 5x. @Princer_Jones

OpenStudy (anonymous):

3x − 4(2x − 5) =3x-4(2x)-4(-5)=3x-8x+20=-5x+20=20-5x Hence true

OpenStudy (anonymous):

@geny55